OpenStudy (aravindg):

The exponent of 7 in 100C50 is??

OpenStudy (experimentx):

??

OpenStudy (anonymous):

Doesn't makes any sense, please do the necessary edit.

OpenStudy (anonymous):

its 1.0089*10^36

OpenStudy (aravindg):

\[\left(\begin{matrix}100 \\ 50\end{matrix}\right)\]

OpenStudy (aravindg):

hw to get tjhis?

OpenStudy (anonymous):

100891344545564193334812497256

OpenStudy (aravindg):

hw do u get that

OpenStudy (aravindg):

i mean can u say the logic

OpenStudy (aravindg):

u see my options ar a)0 b)2 c)4 d) none of these i need to substantiate my answer

OpenStudy (experimentx):

LOL ... what the hell was question?

OpenStudy (aravindg):

what is exponent of 7 in \[\left(\begin{matrix}100 \\ 50\end{matrix}\right)\]

OpenStudy (aravindg):

@FoolForMath

OpenStudy (anonymous):

you can calculate 100c50 which is 1.0089*10^29 and give the power 7 answer is 1.0089*10^36

OpenStudy (aravindg):

i cant use calculator

OpenStudy (experimentx):

exponent of 7 ?? any example of such kind ... though i am smart ... i am really thick headed.

OpenStudy (anonymous):

There is something wrong with the question.

OpenStudy (aravindg):

bt it came in my test :(

OpenStudy (anonymous):

Maybe its asking, "what is the highest power of 7 that divides 100C50"? basically, if you had to prime factor 100C50, what would be the power of 7

OpenStudy (anonymous):

lemme see take 100 move the c divide the possibility of 50 and that leaves me with......I GOT NOT CLUE

OpenStudy (aravindg):

hmm..any logic explanation anyone?

OpenStudy (experimentx):

english must be arch enemy of mathematics

OpenStudy (anonymous):

im illogical jim..

OpenStudy (aravindg):

i cant use calculator thats the main prob

OpenStudy (aravindg):

so i need logical reason

OpenStudy (anonymous):

if the question is what is the power of 7 when you prime factor 100C50, its 0. 7 doesnt divide it.

OpenStudy (aravindg):

thx evryone i will post nxt qn hel p me there tooo

OpenStudy (experimentx):

there's answer in google

OpenStudy (aravindg):

really? giv me the link

OpenStudy (ash2326):

We have \[\frac{100!}{50!{50!}}\] so it's \[\large \frac{(100*99.....51)}{50!}\] let's only write multiples of 7 \[\large \frac{(98.91.84.77.70.63.56)}{7.14.21.28.35.42.49}\] Now we'll write only in terms of powers of 7 \[\large \frac{7^2.7.7.7.7.7.7}{7.7.7.7.7.7.7^2}\] so we end up with \[\huge 7^0\]

OpenStudy (anonymous):

Since:\[\left(\begin{matrix}100 \\ 50\end{matrix}\right)=\frac{100\cdot99\cdot98\cdots 51}{1\cdot 2\cdot 3\cdots 50}\] To find the power of 7 in the numerator, you only need to count the multiples of 7. that gives:\[56, 63, 70, 77, 84, 91, 98\]which would give 7^8 since 98 has 7^2 in it. In the same fashion, the denominator has 7^8 in it. So 7 will cancel out completely.

OpenStudy (aravindg):

so is the answer 0 or none of these?

OpenStudy (ash2326):

we end up with \[\huge 7^0\ \to0\]

OpenStudy (aravindg):

wow thx

OpenStudy (aravindg):

i liked rhis qn

OpenStudy (anonymous):

@ash2326 anything^0 is 1

OpenStudy (ash2326):

I meant that the exponent of 7 is 0 :D

OpenStudy (anonymous):

he knows that, hes saying that there are 0 7's in 100C50

OpenStudy (aravindg):

hehe

OpenStudy (aravindg):

k nexxt qn get ready

OpenStudy (anonymous):

have fun, i need to be working on a take home exam.